The question on my homework is Find the perpendicular line thru the coordinate given

How do you graph 3x+4y=12 if the coordinate given to you is(5,-2) and the line has to be perpindicular?

Answer by stanbon(75887)

How do you graph 3x+4y=12 if the coordinate given to you is (5,-2) and the line has to be perpendicular?
---------------------
Putting the given line into slope-intercept form you get the following:
y=(-3/4)x+3.
That line has a slope of -3/4.
-------------
The slope of a line perpendicular to the given line will have
a slope of (4/3).
----------
The line you want will have form y=mx+b.
You are told that y=-2 when x=5.
So, -2=(4/3)5+b
-2=20/3+b
-6/3=20/3+b
b=-26/3
EQUATION:
y=(4/3)x-(26/3)
Cheers,
Stan H.